Back to the NIKA2 main page

Daily Reports

Because of the pandemic, no observers can come to the telescope. This pool is entirely in remote mode.

Nov 17th, Tuesday

16h00 - 21h00; Bilal: System check, acquisition started, observing setup. TAPAS is not recording new scans (reported). Opacity is pretty high (0.7 < tau_225Ghz < 1, strong instability). Monitoring strong pointing sources aiming for reasonable point corrections. Science observations started: project 029-20.

21h00 - 24h00; Laurence and Frédéric: We take over from Bilal and continue observing on project 029-20 until the targets reached El.~20 deg.

Nov 18th, Wednesday

00h00 - 01h00; Laurence and Frédéric: Opacity increases up to 0.6-0.7, with large fluctuations. Targets of NIKA2 Pool scheduled projects are not feasible.

01h00 - 09h00; Ángel; Stavroula and Suzanne since early morning: taking over at ~01h and switching to EMIR; observing the four B213* sources (3 mm) of the EMIR backup project 011-20. Observations of the EMIR backup project 078-20 started.

09h00 - 15h00; Mónica, Stavroula, Suzanne, Carsten: observations for EMIR project 078-20 continued. Opacity going down; its mean value dropped up to about 0.45 around noon. Bilal helps with switching over to NIKA2 at 13h00 to start with 029-20.

15h00 - 21h00; Bilal: take over and continuing with NIKA2. Observing for project 029-20 since then.

21h00 - 24h00; Laurence and Emmanuel (Artis): Observations for project 029-20 continued. Opacities better than expected from the forecasts, with a nice opacity window from 21:00 and 00:00UT with a mean opacity around 0.3, that allows us to complete a whole iteration on sources. Note that the last 30 min spent on the project sources reached low elevations, from 21 to 14 degrees

Nov 19th, Thursday

00h00 - 03h00; Laurence and Emmanuel (Artis): We went to Uranus for pointing/focus and short calibration scans (calib_1scan). Opacity started rising soon after 00:00 UT. Strong indications of atmospheric instabilities from tau225 fluctuations and beam profiles on pointing scans between 00:30 and 1:30. We did two skydips, hoping they could be used despite instabilities. Indication of improvements after 1:30. We launched 2 pointing "probe" scans at high elevation and found reasonable beam profiles.

03h00 - 09h00; Alexandre: Mean opacity between 0.3 and 0.4 with some spikes during the shift. Two repetitions were carried out on TMC1-C and TMC1-CP for project 006-20. One repetition on JINGLE0 for project 061-20 after that.

09h00 - 15h00; Stavroula, Suzanne, Carsten: Weather stable (tau~0.3). We got one repetition (8 scans) on JINGLE16 of project 061-20. Continued since 09:30 on NGC2976 (24 scans) down to 30deg elevations and then on NGC5194 (2 scans) (project 160-16). TAPAS and (as no IMBFits) piic/ql frozen for a while (IDL/rta used instead). So, scans 61-66 haven't appeared yet in the TAPAS archive. At 13:54 local time the telescope decided to move to the zenith for no obvious reasons. The operator intervened and it did not happen again.

15h00 - 21h00; Bilal: Continuing on NGC5194 of project 160-16. Stopped at 18:30 due to wind gusts.

21h00 - 24h00; Frédéric, Juan: Observations resumed at ~22h. Weather conditions stable with tau just above 0.3. Pointing and focus on Uranus when starting the session. 12/16 scans on MOOJ2206, and 1/2 series of 16 scans MOOJ0319 (project 077-19).

Nov 20th, Friday

00h00 - 03h00; Frédéric, Juan: 2/2 series of 16 scans on MOOJ0319. Beammap on Uranus: 20201120s23.

03h00 - 09h00; Longji: Stable conditions, with 0.2 < tau < 0.3. Started with 1 series of 16 scans on MOOJ0319 (077-19). 12 scans on COSMOS field (192-16) from 04h until the end of the shift.

09h00 - 15h00; Stavroula, Suzanne, Carsten: We observed 2 galaxies of the project 160-16 (8 scans of NGC2976 and 4 scans of NGC5194). As far as we know, we completed the target NGC2976, but we are going to check the data to ensure it. The weather conditions were cooperated (with tau~0.2) and pretty stable for most of the shift. Telescope got stuck 13:15 UT.

15h00 - 21h00; Ankur, Ángel: Quiet stable conditions, 0.2 < tau < 0.4 during the shift with smooth variations. We got 3 repetitions on MOOJ0226 (project 077-20). The first repetition on ACTJ0119 of project 199-20 was started during the final part of the shift.

21h00 - 24h00; Frédéric+Florian: Stable conditions, 0.2 < tau < 0.3. Ran 6 repetitions on ACTJ0119 and 4 repetition for ACTJ0240 (project 199-16). We made a beammap (Uranus, el=66 deg, tau=0.25): 20201120s235.

Nov 21st, Saturday

00h00 - 03h00; Fréderic+Florian: Conditions start to degrade a bit. We moved to 006-20 (NB: the pointing source requested in the README is at a 160 deg azimuth offset from TMC1-*, and at El. < 20 deg, when TMC1-* are > 70 deg).

03h00 - 09h00; Guilaine: Mean opacity between 0.3 and 0.4 with some spikes. 2 hours on TMC1-C. Focus & pointing very stable. 1.5 hours on TMC1-CP. Tau very stable and around 0.34 for the last hour. Then switch to 061-20 (because TMC1-CP has too low elevation). Observed JINGLE10 for ~35 minutes then go to JINGLE16 (higher elevation) for another 35 minutes; tau is increasing.

09h00 - 15h00; Jean-François: Opacity a little above 0.4 and dropped to ~0.3 during the shift with some instabilities. Observed one repetition JINGLE16 (061-20). After that, full steam on id34 (075-20): 3 repetitions, with tau mostly OK. Calibrations on MWC349, CRL2688, NGC7027 + focus, and a skydip.

15h00 - 21h00; Bilal: Starting with tau~0.25 and going down. Observations on id34 (075-20) continued. After that, 2 reps. on MOOJ2206 (077-26) were completed. The shift finishes observing ACTJ0119 (199-16) with opacity below 0.15. Quite stable during the shift.

21h00 - 24h00; Florian+Frédéric: We observed ACTJ0119 and ACTJ0240 (199-16) with a very good tau (0.11 stable)

Nov 22nd, Sunday

00h00 - 03h00; Florian+Frédéric: We made a beammap on Uranus (20201122s4, tau=0.11, el = 58 deg), and moved to 076-20. When the source was too low, we went back to 199-16 (ACTJ0240). When this source was too low also, we have done a skydip. Then we moved to 077-20.

03h00 - 09h00; Longji: The observation of 077-20 was finished at 3:30 am (CET). I started the science observation of 192-16 (N2CLS) at around 4 am and stopped at around 9 am when the telescope went out of the shadow of the mountains. The weather was quite stable with tau ~ 0.15. A total of 32 scans (4 repetitions) on GOODS-N were obtained during the shift.

09h00 - 15h00; Jean-François: I observed sources HD143894 and HD150378 of project 039-20 (tau < 0.20; temp about 5°C but humidity <10%, according to Iram monitoring). Atmospheric fluctuations seen with nk_rta reduction were amazingly low. Focus was done during the middle of the run.

15h00 - 21h00; Bilal: Opacity < 0.2; Observations on HD143894 continued. Then, 2 reps. on NGC628 (160-16).

21h00 - 24h00; Juan, Miren, Emmanuel: Pointing focus on Uranus. ACTJ0119 (199-16): 1 rep. with different angles 22.5 67.5 112.5 157.5 and 7 rep with 0 45 90 135.

Nov 23rd, Monday

00h00 - 03h00; Juan, Miren, Emmanuel: Beammap (scan 8) with Uranus at El.~50 deg. One rep. on MOOJ0319 (077-20).

03h00 - 09h00; Alexandre: After observing 037-20 (1 rep.), I switched to 192-16, but I had to stop because the telescope got stuck, and refuse to move. On strike since 06:35 UT.

09h00 - 15h00; Jean-François: Telescope stuck the first hour. Then, focus and GOODS-N (192-16) for one hour. Then, focus and id34 (075-20) for 2.5 hours (2 reps.). After that, I could not focus for about 45 minutes and finally, the telescope did not take any command. I point to the fact that the long observation of id34 was at about 20°-30° from the Sun.

15h00 - 21h00; Bilal: Starting at 16:40 due to telescope failure, and moving on to project 160-16 (NGC0628) until 9pm.

21h00 - 24h00; Juan, Emmanuel, Miren: focus and pointing on Uranus. Calibration scan on Uranus. 8 repetitions on ACTJ0119 (199-16). Then, a skydip.

Nov 24th, Tuesday

00h00 - 03h00; Juan, Emmanuel, Miren: after a focus and a beammap on Uranus, we did two repetitions on ACTJ0240 (199-16). We moved to 077-20 to observe MOOJ0319.

03h00 - 10h30; Matthieu: We started the shift by observing 077-20 during one hour. Then, we performed a skydip. After minor troubles with the focus, we observed COSMOS (192-16) until sunrise and then GOODS-N (192-16) until the technical time (10h30). The weather was excellent.

DailyReportsNika2Pool171120 (last edited 2021-06-22 14:08:15 by NikaBolometer)